Can someone help me on this?
Ching leaves San Francisco and travels toward Los Angeles at 50 mi/h. An hour later, Phil leaves Los Angeles and travels toward San Francisco at 60 mi/h. If the two cities are 380 miles apart, how many hours will it take for Ching to meet Phil?
Thanks, Adriana

Answer by tutorcecilia(2144)

distance = (rate)(time) [Use the formula for distance while traveling (d=rt)]
.
Ching's distance=rt=(50mph)(unknown time)
Phil's distance=rt= 60mph)(unknown time - 1hour)
Ching's distance (rt) + Phil's distance(rt) = 380
So, 50t+60(t-1)=380
50t+60t-60=380
110t=60+380
110t=440
t=4 hours
.
Check by plugging all of the variables back into the original equation:
50t+60(t-1)=380
50(4)+60(4-1)=380
200+180=380
380=380 [Checks out]